### v1

# Mammo-Clustering:A Weakly Supervised Multi-view Global-Local Context Clustering Network for Detection and Classification in Mammography

Abstract
Breast cancer has long posed a significant threat to women’s health, making early screening crucial for mitigating its impact. However, mammography, the preferred method for early screening, faces limitations such as the burden of double reading by radiologists, challenges in widespread adoption in remote and underdeveloped areas, and obstacles in intelligent early screening development due to data constraints. To address these challenges, we propose a weakly supervised multi-view mammography early screening model for breast cancer based on context clustering. Context clustering, a feature extraction structure that is neither CNN nor transformer, combined with multi-view learning for information complementation, presents a promising approach. The weak supervision design specifically addresses data limitations. Our model achieves state-of-the-art performance with fewer parameters on two public datasets, with an AUC of 0.828 on the Vindr-Mammo dataset and 0.805 on the CBIS-DDSM dataset. Our model shows potential in reducing the burden on doctors and increasing the feasibility of breast cancer screening for women in underdeveloped regions.
Index Terms:
Artificial intelligence, Breast cancer, Deep Learning, Mammography, Medical imaging.
I Introduction
Breast cancer remains the most prevalent malignant tumor among women worldwide[1]. In today’s global context, annually, more than several million women are diagnosed with breast cancer, which constitutes approximately 25% of all cancer cases among women [2]. Recent studies have highlighted that breast cancer has surpassed cardiovascular diseases as a leading cause of premature mortality globally[3] [4]. However, breast cancer is also one of the malignancies for which prevention and treatment strategies are clearly effective and efficacious[5].
Early detection is vital for reducing breast cancer mortality rates [6] [7]. Early detection can also adopt less aggressive treatment plans to improve breast retention rates and reduce the physical and psychological burden on patients. Early detection can also allow for less aggressive treatment options, reducing the physical and psychological burden on patients [8]. Additionally, it can help in identifying high-risk individuals who may benefit from preventive measures, ultimately contributing to improved overall public health outcomes.
Mammography, a low-dose X-ray technique[9], is crucial for early breast cancer detection, improving breast-conserving treatment rates and reducing mortality by over 20% [10]. It effectively identifies tumors too small to be felt, facilitating early intervention. Studies show significant mortality reduction in women aged 50 to 74 after 7 to 9 years, regardless of screening frequency [11]. Its non-invasive nature and detailed imaging enhance diagnosis and treatment planning.
However, mammography is not without its limitations, particularly concerning the risk of misdiagnosis [12]. Double reading has been proposed as a solution to reduce missed diagnoses, yet this approach significantly increases the workload for radiologists [13]. Given that the majority of mammography results are normal, the repetitive and resource-intensive nature of double reading poses a substantial burden on healthcare systems [14] [15]. Furthermore, the high costs and resource demands associated with traditional radiologist-led mammography screenings often restrict access to these services in less developed regions [16].In resource-limited settings, it is crucial to identify which interventions are most effective and feasible in reducing overall breast cancer mortality. Some researchers are dedicated to exploring hybrid screening methods that are more suitable for women in impoverished developing countries [17]. Additionally, related studies have proposed the concept of mobile mammography services [18].
In response to these challenges, substantial efforts have been made to incorporate computer-assisted detection systems to alleviate the burden on radiologists [19]. The integration of artificial intelligence (AI) for autonomous breast cancer prevention and early detection is becoming increasingly mainstream.
Relevant studies have initiated the development of breast cancer detection and classification models based on deep transfer learning [12].
Some studies have attempted to integrate digital mammography and digital breast tomosynthesis in order to detect suspected cancerous regions on mammography images [20].
However, some researchers recently collected mammographic images obtained from different breast X-ray imaging systems and investigated their performance in intelligent early breast cancer screening tasks. They found that the effectiveness of early breast cancer screening systems is often significantly constrained by the quality, accessibility, and standardization of mammographic imaging data[21] [22] [23]. Addressing these challenges is crucial for enhancing the efficiency and accessibility of breast cancer screening across diverse populations. In addition, the inherent complexity and high resolution of mammography images pose substantial challenges for AI algorithms. These images contain intricate anatomical details that are difficult for neural networks to comprehensively extract and integrate into a unified representation.
Traditional approaches to breast cancer detection and classification from mammography images predominantly rely on supervised learning techniques[24] [25] [26]. These methods typically employ object detection networks to extract feature information from mammography images, followed by a series of operations such as feature fusion, prediction, localization, and classification.
Due to the influence of mammography’s feature complexity, and breast cancer detection and classification are constrained by the quality of object detection networks, often resulting in issues such as missed detections and bounding box deviations. Moreover, the scarcity of annotated data poses significant challenges to traditional supervised learning methods in this field.
In response to these limitations, the GMIC (Global and Local Mammography Image Classification) [27] architecture introduces a novel approach based on weakly supervised learning for mammography detection and classification. This framework circumvents the constraints imposed by data scarcity, offering a promising alternative to traditional methodologies.
In traditional machine learning, each sample has a distinct label, but in multiple instance learning (MIL), samples are grouped into “bags” with known labels, while individual “instance” labels are unknown. A bag is labeled positive if at least one instance is positive; otherwise, it is negative [28]. MIL enhances generalization by training on diverse examples, reduces overfitting by exposing models to various scenarios, and improves prediction accuracy by aggregating information across instances. It is also effective for learning from weakly labeled data, useful when precise annotations are costly or difficult to obtain [29] [30].
In mammographic images, an image can be considered a “bag” containing multiple “instances,” which are different image regions or patches. Multi-Instance Learning (MIL) methods allow training with bag-level labels without precise annotations. Even if we do not know which specific instance is the lesion area, the entire bag is labeled as positive if at least one instance within it is positive [31] [27]. We can use this approach to avoid being constrained by the complexity of dataset construction. Our model draws on the concept of multi-instance learning, specifically focusing on selecting important patches and features.
Multi-view learning is a machine learning paradigm that utilizes multiple distinct sets of features, or “views,” to improve learning performance. Each view provides complementary information about the data, enhancing the model’s ability to understand complex patterns [32] [33]. Multi-view learning is employed in various domains such as image and video analysis, natural language processing, and bioinformatics. It is particularly useful in scenarios where data can be naturally divided into different perspectives [34] [35]. The primary advantages of multi-view learning include improved generalization performance, robustness to noise, and the ability to leverage complementary information from different views. By integrating diverse perspectives, multi-view learning can often achieve better accuracy and more insightful models compared to single-view approaches [32] [33]. The relevant study introduces the concept of multi-view learning into the classical generalized eigenvalue proximal support vector machine (GEPSVM) to demonstrate that multi-view learning can enhance model performance by coordinating the complementarity and consistency between different views [36]. The challenges of multi-view learning include effectively integrating information from different perspectives, as data from various viewpoints may require alignment or synchronization, which can be complex in practical applications, particularly when some perspectives may contain noise or misleading information [33].
Given that mammographic images are typically captured from different angles, such as craniocaudal (CC) and mediolateral oblique (MLO) views, and that the symmetry between the left and right breasts in the same view is considered a criterion for health in clinical diagnosis, a multi-view learning strategy is often employed. This approach leverages the complementary information from these different angles to enhance classification performance [37][38].
Convolutional networks, as the most common feature extraction paradigm, are renowned for their ability to capture spatial hierarchies and local patterns in data, making them highly popular in image processing tasks [39] [40]. However, recently, Vision Transformers (ViT) based on attention mechanisms have posed a challenge to convolutional networks [41]. By employing global self-attention operations to adaptively integrate information from patches, they have achieved impressive results.
Recent work addressing the shortcomings of these two feature extraction paradigms has significantly advanced the field of computer vision. Individuals often divide into two factions, each supporting either the convolutional or the Vision Transformer (ViT) paradigm, and focusing on their respective optimizations. Some studies have demonstrated that ResNet, with appropriate training schemes and minor modifications, can perform on par with ViT [42]. The relevant study employs an asymmetric encoder-decoder architecture to randomly mask patches of the input image and reconstruct the missing pixels, enabling efficient and effective training of large models [43]. As the competition between these two paradigms evolves, there is also a continuous integration of both [44]. Recent research, inspired by Vision Transformers (ViT), has achieved more powerful models by replacing a series of small kernels in convolution with several large convolutional kernels.
Nonetheless, beyond convolution and attention mechanisms, there are more possibilities in feature extraction tasks. The success of graph networks has also demonstrated additional potential [45]. Related work employs an anatomy-aware graph convolutional network (AGN) for early breast cancer screening in mammography, achieving promising results [46]. Recent related work directly considers the use of MLP layers for spatial interactions in the representation of image features[47].
Although clustering, as a traditional feature extraction paradigm, has gradually faded from prominence with the development of deep learning [48]. The relevant study segments the image into multiple regions by grouping a set of pixels with common characteristics. The sparsity and simplicity required for clustering demonstrate its satisfactory inter-pretability[49]. Additionally, it can connect representations of point clouds and images, reflecting its remarkable generalization capability .
We propose a Weakly Supervised Multi-instance Multi-view Mammography Classification Network, which effectively utilizes unlabeled or partially labeled data for breast cancer detection and classification.
  • $\bullet$
    Our work is the first to apply a non-CNN and non-attention mechanism image feature extraction method, namely Context Clustering, to the early screening task of mammography for breast cancer.
  • $\bullet$
    Our work achieves state-of-the-art accuracy in the early screening task of breast cancer mammography while maintaining the lowest number of parameters.
  • $\bullet$
    We propose a novel fusion mechanism that integrates global information, feature-based local information, and patch-based local information, placing greater emphasis on local details compared to previous methods.
II Method
Our research aims to develop a rapid and reliable breast cancer early screening system to alleviate the burden on medical professionals and provide screening opportunities for women’s health in underdeveloped regions.
In this section, we describe the weakly supervised multi-instance multi-view network architecture we propose. In the first subsection, we formulate the steps involved in early breast cancer screening through mammography. In the second subsection, we provide a detailed description of the framework we propose. In the third subsection, we provide a detailed introduction and evaluation of the dataset used. In the four subsection, we introduce the evaluation metrics used to assess the effectiveness of early breast cancer screening.
II-A Overall Framework
Figure 1: Architecture of the proposed model. Images from four perspectives are enhanced into point sets and processed via a multi-level Context Clustering module, Global Coc, to extract global information. This module includes point reducers and Context Cluster Blocks. The ROISelectModel utilizes this global information to select patch-based images, which are processed through another Context Clustering module, Local Coc, to generate patch-based local information. This is fused with feature-based local information derived from the Global information to produce Local information. Subsequently, Local and Global information are combined to create single-view fusion information. Fusion information from each perspective is integrated across views and regressed to produce the final output.
The proposed model for mammography classification can be formulated as follows:
For each image $I$ in the given view, we enhance all points into 5-dimensional information points containing color and position data to obtain the set of points $S\in\mathbb{R}^{5\times w\times h}$, where $w\times h$ is equal with the number of points.
II-A1 Global Information Extraction
The set $S$ is input to the first point clustering network to obtain global information $F_{g}$ and Saliency-map $I_{map}$:
$F_{g},I_{map}=f_{\textbf{global}}(S)$ (1)
Where $f_{\textbf{global}}$ represents the point clustering network for global information extraction. $F_{g}$ is extracted global information from global network. The saliency-map extracted by the global network, denoted as $I_{map}$.
$P=f_{roi}(I_{map})$
The $P$ is a set of location information, representing the positions of n example patches selected by the ROI selection function $f_{roi}$:
$P=\{p_{1},p_{2},\ldots,p_{n}\}$
where $p_{n}$ is a coordinate representing a position, written as $(x_{n},y_{n})$.
With $p_{n}$, we can extract n patches $\tilde{I}$ from the original image $I_{i}$ and extract n feature-based local information $F_{fl}$ from the global feature information $F_{g}$.
$F_{fl}^{n}=F_{g}(x_{n},y_{n})$ (2)
where $F_{fl}^{n}$ is means n-th feature-based local information.
II-A2 Local Information Extraction
Each selected patch $\tilde{I}_{n}$ is treated as a new image, re-enhanced based on each point to obtain its five dimensional point set $\tilde{S}_{n}$. And processed through a second point clustering network $f_{\textbf{local}}$ to obtain patch-based local information $F_{pl}^{n}$ :
$F_{pl}^{n}=f_{\textbf{local}}(\tilde{S}_{n})$ (3)
where $f_{\textbf{local}}$ represents the point clustering network for local information extraction, similar to global clustering network $f_{\textbf{global}}$.
II-A3 Information Fusion and Attention Mechanism
The local information $F_{l}$ from all patches is fused with all feature-based local information and all patch-based local information :
$F_{l}=F_{fl}\oplus F_{pl}$ (4)
where Operation $\oplus$ overlays two features.
after that processed through an attention mechanism to enhance relevant features:
$F_{a}=f_{attention}(F_{l})$ (5)
Then, the attention-enhanced information is fused with the original global information, resulting in multi-instance fusion information from single-view:
$F_{f}=f_{\textbf{fuse}}(F_{g},F_{a})$ (6)
II-A4 View Fusion and Classification
Process the images $I_{view}$ from the four views (bilateral craniocaudal (CC) and mediolateral oblique (MLO)) using the aforementioned procedure to obtain single-view fusion information. This information is then integrated for multi-view fusion, which is used for the final binary classification, resulting in the early screening model’s output:
$F_{fusion}={f}^{\prime}_{fuse}(F_{f}^{\text{lcc}},F_{f}^{\text{lmlo}},F_{f}^{\text{rcc}},F_{f}^{\text{rmlo}})$ (7)
where $F_{f}^{\text{lcc}}$ represents the fusion feature of LCC images, other similar situations. ${f}^{\prime}_{fuse}$ is another fusion structure different from $f_{\textbf{fuse}}$. In addition to integrating features $F_{fusion}$, ${f}^{\prime}_{fuse}$ also merges $F_{g}$ and $F_{l}$ from different views.
$F_{global}={f}^{\prime}_{fuse}(F_{g}^{\text{lcc}},F_{g}^{\text{lmlo}},F_{g}^{\text{rcc}},F_{g}^{\text{rmlo}})$
$F_{local}={f}^{\prime}_{fuse}(F_{l}^{\text{lcc}},F_{l}^{\text{lmlo}},F_{l}^{\text{rcc}},F_{l}^{\text{rmlo}})$
Finally, $F_{fusion}$ performs the final classification through $f_{cls}$.
$\text{res}=f_{cls}(F_{fusion})$ (8)
This formulation encapsulates the entire process of the weakly supervised multi-instance multi-view network for mammography classification.
II-B Detailed Network Structure
II-B1 From Image to Set of Points
The scale of an image can be expressed as (3, h, w), where 3 represents the RGB channels, and h and w are its height and width. We enhance each pixel by considering it as a 5-dimensional data point containing color and positional information (r, g, b, x, y). After this enhancement, the image can be represented as a set of $h\times w$ 5-dimensional data points, with a scale of $(h\times w,5)$. We can then perform feature extraction through simple clustering. From a global perspective, the image is viewed as a collection of unordered discrete data points with color and positional information. Through clustering, all points are grouped into clusters, each containing a centroid. Since each point in the set includes color and positional information, this clustering implicitly incorporates spatial and image information.
II-B2 Context Cluster Block
We employ context cluster blocks for hierarchical feature extraction, a paradigm similar to convolutional networks. At the beginning of each stage, we utilize a point reducer to decrease the number of points, thereby enhancing computational efficiency. Subsequently, a series of context cluster blocks are used to extract deep features and adaptively assign aggregated features to each uniformly selected anchor point within the cluster based on similarity, and connect and fuse the nearly points through linear projection. Finally, we perform a point-wise averaging operation on the output of the last layer.
Figure 2: The figure illustrates the clustering visualization following five steps: selecting central anchor points, identifying neighbors for each anchor, calculating features for each anchor, performing similarity analysis based on these anchors, and representing all clusters on the chart.
As illustrated in Figure 2, the input image undergoes point set transformation, and then, in step a, $n$ central anchor points are uniformly selected in the space. the method is similar to those in SuperPixel [50] and SLIC [51].
The selected central anchor points are highlighted with red boxes in the figure. In step b, for each central anchor point, $k$ neighbors are identified, indicated by arrows in the figure. The value of $k$ can be 4 or 8, as determined manually, and it can also be the four neighbors in the up, down, left, and right directions, in which case $k$ equals 4.
Step c involves calculating the features of a central anchor point determined by itself and its $k$ neighboring points, illustrated in the figure for the case where the number of neighbors is 8. The calculation process is:
$P^{x}=\frac{\left(P^{x}+\sum_{q=1}^{k}p_{q}^{x}\right)}{k+1}$
where $P^{x}$ represents a $x$ dimension of the central anchor point, $x\in\{r,g,b,h,w\}$. And $p_{q}^{x}$ represents the $n$-th neighbor point in the $x$ dimension, $q\in\{0,1,2,\ldots,k\}$.
After computing the features for all central anchor points, a similarity analysis is conducted between all points in the point set and each central anchor point’s features in step d. Each point is assigned to the cluster of the central anchor point with which it has the highest similarity. The steps for the similarity analysis is conducted by computing the pairwise cosine similarity matrix between a point and set of central points:
$f(P_{i},P^{\prime}_{j})=\frac{P_{i}\cdot P^{\prime}_{j}}{|P_{i}||P^{\prime}_{j}|}$
where $f(P_{i},P^{\prime}_{j})$ is pairwise cosine similarity compute. $P_{i}$ is $i$-th central anchor point, $i\in\{0,1,2,\ldots,n\}$. $P^{\prime}_{j}$ is $j$-th point in the image point set.
Finally, in step e, all clusters are combined, resulting in the desired clustering outcome for the entire image.
Since each point contains both feature and positional information, the calculation implicitly emphasizes the point’s distance (locality) and feature similarity. Each point is then assigned to the most similar center, resulting in c clusters. During this process, clusters may contain varying numbers of points. In extreme cases, some clusters may contain zero points, rendering them redundant.
II-B3 ROI Selection Module
Firstly, the dimensions of the region of interest mapping are required to be consistent with the dimensions of the saliency-map.
$h_{crop}=h_{patch}\times h_{map}\div h_{I}$
$w_{crop}=w_{patch}\times w_{map}\div w_{I}$
where $h_{patch}$ and $w_{patch}$ are user-defined for the patch image size. $h_{map}$ and $w_{map}$ are the size of the saliency-map. $h_{I}$ and $w_{I}$ are the size of the original image data. $h_{crop}$ and $w_{crop}$ are dynamically adjusted based on the original image and saliency map scales. Subsequently, the saliency-map is normalized and divided into regions with a height of $h_{crop}$ and a width of $w_{crop}$ regions for greedy ROI search. In each iteration, the algorithm greedily identifies each region and selects the one with the largest total weight, as determined by average pooling, among all current regions. The coordinates of this region are added to a list, and a mask flag is applied to the region to prevent redundant selection. The coordinates of these regions will be mapped to the size of the original image to obtain patch-based images.
Figure 3 visualizes the patch-based images selected by the ROISelectModule, along with the patches’ positions on the source image and their comparison with the locations of suspicious lesions.
Vindr-Mammo CBIS-DDSM
AUC ACC F1 score AUC ACC F1 score Params
Single-View Res[37] $0.727\pm 0.02$ 0.783 0.619 $0.719\pm 0.02$ 0.591 0.558 1477025
Single-View Swin-transformer $0.731\pm 0.02$ 0.651 0.594 $0.724\pm 0.02$ 0.601 0.599 14184625
GMIC[27] $0.793\pm 0.02$ 0.624 0.518 $0.778\pm 0.02$ 0.712 0.705 22487298
Multi-View res[37] $0.740\pm 0.02$ 0.753 0.567 $0.731\pm 0.02$ 0.676 0.630 6128546
MaMVT[38] $0.770\pm 0.02$ 0.918 0.647 $0.749\pm 0.02$ 0.649 0.649 30730082
Multi-View GMIC[52] $0.797\pm 0.02$ 0.637 0.521 $0.781\pm 0.02$ 0.719 0.699 22686871
Mammo-Clustering(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.694 $\textbf{0.805}\pm 0.02$ 0.709 0.709 9805459
TABLE I: Performance of each model on two datasets
II-B4 Attention Module
In our model, we use two different attention mechanisms to separately fuse multi-instance and multi-view information.
Similar Multi-instances information fusion: After the global network, the ROI select module to choose $k$ patch-based images, a number set manually. This implies not all patch-based images carry beneficial information, and some may be redundant. Considering and integrating all the information from these patch-based images could significantly impair our network. Therefore, an attention module is added before integrating local and global information, allowing the model to learn how to filter out irrelevant local information.
The attention mechanism receives feature representation of patch-based images $F_{l}$, shaped as $(batchsize,k,dim)$, where $batchsize$ and $k$ are manually set parameters; the former defines the batch size during training, while the latter specifies the number of patches required for multiple instance learning. The size of $dim$ varies depending on the model.
First, we use a neural network layer with simple linear transformations $f_{weights}$ and softmax function to compute attention weights $W_{att}$.
$W_{att}=\text{softmax}(f_{weights}(F_{l}))$
$F_{a}=F_{l}\odot W_{att}$
where $\odot$ represents the stationary point multiplication algorithm. Subsequently, the attention weights $W_{att}$ are multiplied pointwise with the feature representation of patch-based images $F_{l}$ is performed to obtain the final implicit representation $F_{a}$.
Multi-view information fusion: In multi-view learning, not all information from each view is necessarily classified as malignant. However, If a single view exhibits malignant characteristics, the instance should be classified as malignant. Therefore, we introduce an attention mechanism to enable the model to autonomously filter out irrelevant view information, enhancing classification accuracy.
The attention mechanism processing is largely consistent with multi-instance fusion attention. However, in multi-view attention, this attention module processes not only the $F_{a}$ fused by the multi-instance attention module but also $F_{g}$ and $F_{l}$. This is because all three features are considered in the loss function for loss calculation.
II-B5 Embedding Module
The embedding module in the model primarily aligns feature-based local information $F_{fl}$ with patch-based local information $F_{pl}$ before their integration. Here, we employ a trainable MLP to align the scales. We designed relevant ablation experiments to verify its effectiveness.
II-B6 Maxpooling Module
We employ max-pooling to fold and align global information $F_{g}$, facilitating better integration with local information fused through the attention module.
II-C Loss Function
We chose a composite loss function to achieve targeted optimization of different components.
$L_{total}=\alpha\cdot L_{global}+\beta\cdot L_{local}+\gamma\cdot L_{fusion}+\delta\cdot L_{map}$
After the multi-view fusion module, we retain not only the fused information for regression but also intermediate features such as global information, local information, and saliency maps. These features are used to compute a composite loss function for precise optimization of each part of the network.And we determine the sensitivity of the loss function to different types of data through component analysis.
$L_{global}$ is calculated using the global information obtained from multi-view fusion and the ground truth values. The loss function chosen here is BCELoss. And $L_{map}$ will be calculated from the saliency-map, it is the weighted average intensity of the saliency-map under the L1 norm. The $L_{global}$, combined with $L_{map}$, indicates the quality of the Global Network and further to determining the adjustment magnitude for the Global Network to enhance the accuracy of locating patch-based images. BCEWithLogitsLoss function is used for both $L_{local}$ and $L_{fusion}$. $L_{local}$ represents the quality of the local network, calculated from local information and ground truth values, determining the adjustment magnitude for the Local Network to enhance the feature extraction capability of the Local Network. $L_{fusion}$ represents the model’s final classification error, driving the overall model training. The weights $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the proportion of each loss, all manually set.
Figure 3: Figures a to d provide a visualization of patch-based images extracted by the model. The green box on the mammography indicates the location of the suspicious lesion, while the blue box represents the patch-based images selected by the model. We can observe that the model’s extracted patch-based images perform exceptionally well, and the magnified images clearly show calcifications and masses.
III Experiment and Result
III-A Datasets
III-A1 Vindr-Mammo
The Vindr-Mammo [53] dataset is a large-scale, annotated collection of digital mammographic images aimed at advancing breast cancer detection and diagnosis through machine learning. It includes thousands of images sourced from diverse populations, with detailed annotations such as lesion types, BI-RADS categories, and precise lesion locations. This dataset is designed to support the development of robust AI models by providing a wide variety of cases, including both normal and abnormal findings, thus enhancing the generalizability and accuracy of diagnostic algorithms.
III-A2 CBIS-DDSM
The CBIS-DDSM (Curated Breast Imaging Subset of the Digital Database for Screening Mammography) [54] dataset is a widely used resource in the field of breast cancer research. It comprises digitized film mammograms, which have been meticulously annotated with information such as lesion boundaries, types (e.g., calcifications, masses), and pathology-confirmed labels (benign or malignant). The dataset also includes patient metadata and additional clinical information, making it an invaluable tool for training and validating computer-aided detection and diagnosis systems. Its comprehensive nature and established use in the research community make it a benchmark for evaluating the performance of mammography-based AI models.
Vindr-Mammo CBIS-DDSM
benign malignant total benign malignant total
train 3614 385 3999 629 660 1289
test 904 96 1000 185 146 331
total 4518 481 4999 814 806 1620
TABLE II: The composition of data for the two datasets.
III-B Evaluating Indicator
In breast cancer early screening models, several evaluation metrics are commonly used to assess the performance of the classification models. Here are the definitions and significance of each metric along with their respective formulas:
III-B1 AUC (Area Under the Curve)
AUC means the area under the receiver operating characteristic (ROC) curve. The ROC curve uses the true positive rate for mammography benign-malignant classification as the y-axis and the false positive rate as the x-axis. It provides an aggregate measure of performance across all possible classification thresholds. A higher AUC value indicates a better model performance, with 1 representing a perfect model and 0.5 a random guess.
[…]
where $TPR(t)$ is the true positive rate at threshold t, and $FPR(t)$ is the false positive rate at threshold t
III-B2 ACC (Accuracy)
Accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined. It gives a straightforward measure of how often the classifier is correct.
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$
where $TP$, $TN$, $FP$ and $FN$ represent the numbers of true positives, true negatives, false positives, and false negatives, respectively.
III-B3 F1 Score
The F1 Score is the weighted average of Precision and Recall. This score takes both false positives and false negatives into account. It is particularly useful when the class distribution is uneven.
[…]
The F1 score reaches its best value at 1 (perfect precision and recall) and worst at 0.
These metrics collectively provide a comprehensive evaluation of the performance of breast cancer screening models, helping to understand their strengths and weaknesses in various aspects of classification.
III-C Comparative Experiment
In this study, we evaluated several models on two datasets: Vindr-Mammo and cbis-ddsm. The performance metrics considered were AUC, ACC, and F1 score.
III-C1 Vindr-Mammo
For the Vindr-mammo, ours model achieved the highest performance across all metrics, with an AUC of $0.828\pm 0.02$, ACC of $0.919$, and F1 score of $0.694$. For single-view task models, the GMIC with a Global-local structure achieves an AUC of 0.793, significantly outperforming ResNet and Swin-Transformer, which had AUCs of $0.727\pm 0.02$ and $0.731\pm 0.02$, respectively. For multi-view task models, our model achieves an AUC of 0.828, significantly surpassing other models. Multi-View GMIC also showed competitive performance with an AUC of $0.797\pm 0.02$, validating the effectiveness of the Global-local architecture, but its ACC and F1 score were lower at $0.637$ and $0.512$.
III-C2 cbis-ddsm
On the cbis-ddsm dataset, the ours model again demonstrated superior performance with an AUC of $0.805\pm 0.02$, ACC of $0.709$, and F1 score of $0.709$. Multi-View GMIC, which had an AUC of $0.781\pm 0.02$ and an F1 score of $0.699$, obtained the highest ACC on this dataset. In this dataset, the advantages of the Global-Local architecture are more pronounced, with ResNet and Swin-Transformer showing significant disadvantages in AUC.
III-C3 Model Complexity
In terms of model complexity, measured by the number of parameters, the ours model had 98.05 million parameters. Other smaller networks often cannot achieve the accuracy of our model and show a significant gap. This is efficient compared to the MaMVT with 30.73 million parameters and the Multi-View GMIC with 22.68 million parameters, considering the performance gains achieved.
III-C4 ROC curve
Figure 4: Figures a and b compare the ROC curves of our model with other Single-view and Multi-view architectures on the Vindr-mammo dataset. Figures c and d present the ROC curves comparison on the CBIS-DDSM dataset.
The ROC curve in figure 4 provides insights that cannot be obtained from tables alone. Analyzing the ROC curve, we observe that most models, except ours, exhibit a concave shape in the middle. This is due to class imbalance in the data, further validating the effectiveness of our model’s architecture.
Overall, our model offers a robust and efficient approach, achieving state-of-the-art performance on both datasets, surpassing the second-best AUC by over 0.02, with fewer parameters. The Global-Local architecture proves effective for both multi-view and single-view models. Additionally, the multi-view learning approach enhances model performance.
III-D Ablation Experiment
III-D1 Different Information Fusion Method
This ablation study aims to demonstrate the effectiveness of our proposed weakly supervised architecture on mammography.
AUC ACC F1 score
Single-view Coc $0.762\pm 0.02$ 0.711 0.627
Multi-View Coc $0.783\pm 0.02$ 0.815 0.658
Ours(mv+gl) $\textbf{0.828}\pm 0.02$ 0.919 0.694
TABLE III: Performance of different information fusion methods on the Vindr-Mammo
The table clearly demonstrates the superiority of our architecture, achieving the highest AUC as well as optimal ACC and F1 scores, indicating its balance in mammography tasks. $mv$ represents a multi-view learning approach, and $gl$ refers to our proposed weakly supervised framework.
III-D2 Different Local Information
We identified two distinct sources of local information: patch-based local information and feature-based local information. Moreover, this feature-based local information has been overlooked in existing work.
This ablation study aims to verify the effectiveness of the mechanism integrating feature-based local information with patch-based local information.
AUC ACC F1 score
patch-based local information $0.810\pm 0.02$ 0.873 0.678
feature-based local information $0.806\pm 0.02$ 0.895 0.659
ours $\textbf{0.828}\pm 0.02$ 0.919 0.694
TABLE IV: Performance of different local information on the Vindr-Mammo
We found that focusing on only one type of local information doesn’t yield better results. The AUC for patch-based local information is 0.810, and for feature-based local information, it’s 0.806. However, combining both achieved the best result, with an AUC of 0.828.
IV Discussion
Figure 5: From figure a to d, the left half of each image shows the original mammogram with annotated suspicious lesions, while the right half presents our contextual clustering visualization, akin to a CNN heatmap and a VIT attention map, with the suspicious lesion locations also outlined. This figure clearly shows that our Context Clustering approach effectively identifies and groups suspicious lesion areas in mammography.
Figure 5 reveals that lesions in mammography predominantly appear in clustered forms. The Coc feature extraction paradigm utilizes clustering of point sets based on color and location information, making it adept at identifying the shape and position of such lesions. This paradigm offers simplicity, resulting in excellent interpretability and generalizability. Therefore, we believe that the Coc network has a strong capability to learn the prior structure of lesions in mammography images.
Single-View GMIC Multi-View GMIC Ours
MDR 0.291 0.277 0.177
TABLE V: The performance of different models in terms of missed detection rate.
We introduce an additional evaluation metric, the missed detection rate (MDR):
$\textit{MDR}=\frac{N_{miss}}{N_{gt}}$
MDR is defined as the percentage of the number of unrecognized suspicious lesion areas $N_{miss}$ relative to the total number of suspicious lesion areas $N_{gt}$. Because, in clinical practice, we are more concerned about lesions being undetected, i.e., false negatives, rather than false positives.
From Table V, we can see that our model achieved the lowest missed detection rate of 0.177, surpassing other model with rates close to 0.1, demonstrating its potential. The other two models, like ours, are weakly supervised methods capable of identifying lesion locations using only classification labels, making them suitable for comparison.
V Conclusions
In this study, we developed a novel weakly supervised multi-view model for early breast cancer screening using mammography images. Unlike conventional feature extraction paradigms such as CNNs and Transformers, our approach employs a context clustering-based method. This strategy allows for the integration of feature-based local information with patch-based local information, enhancing the model’s ability to capture nuanced image details. Furthermore, by incorporating multi-view mammography image features, our model effectively leverages complementary information from different perspectives. This comprehensive approach addresses the limitations of single-view analysis and improves diagnostic accuracy. The model’s performance was rigorously evaluated on two publicly available datasets, Vindr-Mammo and CBIS-DDSM, where it achieved state-of-the-art accuracy, demonstrating its potential as a robust tool for early breast cancer detection.

### v2

@@
-Mammo-Clustering:A Weakly Supervised Multi-view Global-Local Context Clustering Network for Detection and Classification in Mammography
+Mammo-Clustering: A Weakly Supervised Multi-view Tri-level Information Fusion Context Clustering Network for Localization and Classification in Mammography
 Abstract
-Breast cancer has long posed a significant threat to women’s health, making early screening crucial for mitigating its impact.
-However, mammography, the preferred method for early screening, faces limitations such as the burden of double reading by radiologists, challenges in widespread adoption in remote and underdeveloped areas, and obstacles in intelligent early screening development due to data constraints.
-To address these challenges, we propose a weakly supervised multi-view mammography early screening model for breast cancer based on context clustering.
-Context clustering, a feature extraction structure that is neither CNN nor transformer, combined with multi-view learning for information complementation, presents a promising approach.
-The weak supervision design specifically addresses data limitations.
-Our model achieves state-of-the-art performance with fewer parameters on two public datasets, with an AUC of 0.828 on the Vindr-Mammo dataset and 0.805 on the CBIS-DDSM dataset.
-Our model shows potential in reducing the burden on doctors and increasing the feasibility of breast cancer screening for women in underdeveloped regions.
+Breast cancer is a significant global health issue, and the diagnosis of breast imaging has always been challenging.
+Mammography images typically have extremely high resolution, with lesions occupying only a very small area.
+Down-sampling in neural networks can easily lead to the loss of microcalcifications or subtle structures, making it difficult for traditional neural network architectures to address these issues.
+To tackle these challenges, we propose a Context Clustering Network with triple information fusion.
+Firstly, compared to CNNs or transformers, we find that Context clustering methods (1) are more computationally efficient and (2) can more easily associate structural or pathological features, making them suitable for the clinical tasks of mammography.
+Secondly, we propose a triple information fusion mechanism that integrates global information, feature-based local information, and patch-based local information.
+The proposed approach is rigorously evaluated on two public datasets, Vindr-Mammo and CBIS-DDSM, using five independent splits to ensure statistical robustness.
+Our method achieves an AUC of $0.828\pm 0.020$ on Vindr-Mammo and $0.805\pm 0.020$ on CBIS-DDSM, outperforming the next best method by 3.1% and 2.4%, respectively.
+These improvements are statistically significant ($p<0.05$), underscoring the benefits of Context Clustering Network with triple information fusion.
+Overall, our Context Clustering framework demonstrates strong potential as a scalable and cost-effective solution for large-scale mammography screening, enabling more efficient and accurate breast cancer detection.
+Access to our method is available at https://github.com/Sohyu1/Mammo_Clustering.
+Index Terms:
+Artificial intelligence, Breast cancer, Deep Learning, Mammography, Medical imaging.
 Index Terms:
@@
 I Introduction
-Breast cancer remains the most prevalent malignant tumor among women worldwide[1].
-In today’s global context, annually, more than several million women are diagnosed with breast cancer, which constitutes approximately 25% of all cancer cases among women [2].
-Recent studies have highlighted that breast cancer has surpassed cardiovascular diseases as a leading cause of premature mortality globally[3] [4].
-However, breast cancer is also one of the malignancies for which prevention and treatment strategies are clearly effective and efficacious[5].
-Early detection is vital for reducing breast cancer mortality rates [6] [7].
-Early detection can also adopt less aggressive treatment plans to improve breast retention rates and reduce the physical and psychological burden on patients.
-Early detection can also allow for less aggressive treatment options, reducing the physical and psychological burden on patients [8].
-Additionally, it can help in identifying high-risk individuals who may benefit from preventive measures, ultimately contributing to improved overall public health outcomes.
-Mammography, a low-dose X-ray technique[9], is crucial for early breast cancer detection, improving breast-conserving treatment rates and reducing mortality by over 20% [10].
-It effectively identifies tumors too small to be felt, facilitating early intervention.
-Studies show significant mortality reduction in women aged 50 to 74 after 7 to 9 years, regardless of screening frequency [11].
-Its non-invasive nature and detailed imaging enhance diagnosis and treatment planning.
-However, mammography is not without its limitations, particularly concerning the risk of misdiagnosis [12].
-Double reading has been proposed as a solution to reduce missed diagnoses, yet this approach significantly increases the workload for radiologists [13].
-Given that the majority of mammography results are normal, the repetitive and resource-intensive nature of double reading poses a substantial burden on healthcare systems [14] [15].
-Furthermore, the high costs and resource demands associated with traditional radiologist-led mammography screenings often restrict access to these services in less developed regions [16].In resource-limited settings, it is crucial to identify which interventions are most effective and feasible in reducing overall breast cancer mortality.
-Some researchers are dedicated to exploring hybrid screening methods that are more suitable for women in impoverished developing countries [17].
-Additionally, related studies have proposed the concept of mobile mammography services [18].
-In response to these challenges, substantial efforts have been made to incorporate computer-assisted detection systems to alleviate the burden on radiologists [19].
-The integration of artificial intelligence (AI) for autonomous breast cancer prevention and early detection is becoming increasingly mainstream.
-Relevant studies have initiated the development of breast cancer detection and classification models based on deep transfer learning [12].
-Some studies have attempted to integrate digital mammography and digital breast tomosynthesis in order to detect suspected cancerous regions on mammography images [20].
-However, some researchers recently collected mammographic images obtained from different breast X-ray imaging systems and investigated their performance in intelligent early breast cancer screening tasks.
-They found that the effectiveness of early breast cancer screening systems is often significantly constrained by the quality, accessibility, and standardization of mammographic imaging data[21] [22] [23].
-Addressing these challenges is crucial for enhancing the efficiency and accessibility of breast cancer screening across diverse populations.
-In addition, the inherent complexity and high resolution of mammography images pose substantial challenges for AI algorithms.
-These images contain intricate anatomical details that are difficult for neural networks to comprehensively extract and integrate into a unified representation.
-Traditional approaches to breast cancer detection and classification from mammography images predominantly rely on supervised learning techniques[24] [25] [26].
-These methods typically employ object detection networks to extract feature information from mammography images, followed by a series of operations such as feature fusion, prediction, localization, and classification.
-Due to the influence of mammography’s feature complexity, and breast cancer detection and classification are constrained by the quality of object detection networks, often resulting in issues such as missed detections and bounding box deviations.
-Moreover, the scarcity of annotated data poses significant challenges to traditional supervised learning methods in this field.
-In response to these limitations, the GMIC (Global and Local Mammography Image Classification) [27] architecture introduces a novel approach based on weakly supervised learning for mammography detection and classification.
-This framework circumvents the constraints imposed by data scarcity, offering a promising alternative to traditional methodologies.
-In traditional machine learning, each sample has a distinct label, but in multiple instance learning (MIL), samples are grouped into “bags” with known labels, while individual “instance” labels are unknown.
-A bag is labeled positive if at least one instance is positive; otherwise, it is negative [28].
-MIL enhances generalization by training on diverse examples, reduces overfitting by exposing models to various scenarios, and improves prediction accuracy by aggregating information across instances.
-It is also effective for learning from weakly labeled data, useful when precise annotations are costly or difficult to obtain [29] [30].
-In mammographic images, an image can be considered a “bag” containing multiple “instances,” which are different image regions or patches.
-Multi-Instance Learning (MIL) methods allow training with bag-level labels without precise annotations.
-Even if we do not know which specific instance is the lesion area, the entire bag is labeled as positive if at least one instance within it is positive [31] [27].
-We can use this approach to avoid being constrained by the complexity of dataset construction.
-Our model draws on the concept of multi-instance learning, specifically focusing on selecting important patches and features.
-Multi-view learning is a machine learning paradigm that utilizes multiple distinct sets of features, or “views,” to improve learning performance.
-Each view provides complementary information about the data, enhancing the model’s ability to understand complex patterns [32] [33].
-Multi-view learning is employed in various domains such as image and video analysis, natural language processing, and bioinformatics.
-It is particularly useful in scenarios where data can be naturally divided into different perspectives [34] [35].
-The primary advantages of multi-view learning include improved generalization performance, robustness to noise, and the ability to leverage complementary information from different views.
-By integrating diverse perspectives, multi-view learning can often achieve better accuracy and more insightful models compared to single-view approaches [32] [33].
-The relevant study introduces the concept of multi-view learning into the classical generalized eigenvalue proximal support vector machine (GEPSVM) to demonstrate that multi-view learning can enhance model performance by coordinating the complementarity and consistency between different views [36].
-The challenges of multi-view learning include effectively integrating information from different perspectives, as data from various viewpoints may require alignment or synchronization, which can be complex in practical applications, particularly when some perspectives may contain noise or misleading information [33].
-Given that mammographic images are typically captured from different angles, such as craniocaudal (CC) and mediolateral oblique (MLO) views, and that the symmetry between the left and right breasts in the same view is considered a criterion for health in clinical diagnosis, a multi-view learning strategy is often employed.
-This approach leverages the complementary information from these different angles to enhance classification performance [37][38].
-Convolutional networks, as the most common feature extraction paradigm, are renowned for their ability to capture spatial hierarchies and local patterns in data, making them highly popular in image processing tasks [39] [40].
-However, recently, Vision Transformers (ViT) based on attention mechanisms have posed a challenge to convolutional networks [41].
-By employing global self-attention operations to adaptively integrate information from patches, they have achieved impressive results.
-Recent work addressing the shortcomings of these two feature extraction paradigms has significantly advanced the field of computer vision.
-Individuals often divide into two factions, each supporting either the convolutional or the Vision Transformer (ViT) paradigm, and focusing on their respective optimizations.
-Some studies have demonstrated that ResNet, with appropriate training schemes and minor modifications, can perform on par with ViT [42].
-The relevant study employs an asymmetric encoder-decoder architecture to randomly mask patches of the input image and reconstruct the missing pixels, enabling efficient and effective training of large models [43].
-As the competition between these two paradigms evolves, there is also a continuous integration of both [44].
-Recent research, inspired by Vision Transformers (ViT), has achieved more powerful models by replacing a series of small kernels in convolution with several large convolutional kernels.
-Nonetheless, beyond convolution and attention mechanisms, there are more possibilities in feature extraction tasks.
-The success of graph networks has also demonstrated additional potential [45].
-Related work employs an anatomy-aware graph convolutional network (AGN) for early breast cancer screening in mammography, achieving promising results [46].
-Recent related work directly considers the use of MLP layers for spatial interactions in the representation of image features[47].
-Although clustering, as a traditional feature extraction paradigm, has gradually faded from prominence with the development of deep learning [48].
-The relevant study segments the image into multiple regions by grouping a set of pixels with common characteristics.
-The sparsity and simplicity required for clustering demonstrate its satisfactory inter-pretability[49].
-Additionally, it can connect representations of point clouds and images, reflecting its remarkable generalization capability .
-We propose a Weakly Supervised Multi-instance Multi-view Mammography Classification Network, which effectively utilizes unlabeled or partially labeled data for breast cancer detection and classification.
+Recent studies highlight that breast cancer, as the most common malignancy among women, has surpassed cardiovascular diseases to become the leading cause of premature mortality among women worldwide [1] [2].
+However, breast cancer is also notably amenable to effective prevention and treatment strategies [3].
+Early detection of breast cancer is crucial for reducing breast cancer mortality and significantly improving patient prognosis [4] [5].
+It will allows for less invasive and more targeted treatment options, thereby reducing the physical and psychological burden on patients [6].
+Mammography is a low-dose, non-invasive X-ray imaging technique [7] that plays a crucial role in the early detection of breast cancer by identifying tumors too small to be palpated, thereby facilitating timely intervention.
+Several studies have suggested that using mammography for early breast cancer screening can significantly reduce mortality by up to 20% [8].
+One aspect of the specificity of the mammography issue is that, as a multi-view imaging technique, mammograms are typically acquired from the craniocaudal (CC) and mediolateral oblique (MLO) angles of both the left and right breasts.
+From a given perspective, the symmetry between the left and right breasts also serves as a critical diagnostic criterion in clinical practice.
+Consequently, employing a multi-view learning strategy can leverage the complementary information provided by these different imaging angles, thereby enhancing classification performance [9] [10].
+Multi-view learning is a machine learning paradigm that leverages multiple feature sets, or “views,” to improve performance by capturing complementary insights.
+Widely applied in fields like image analysis, NLP, and bioinformatics, it enhances generalization, robustness to noise, and accuracy by integrating diverse perspectives, often outperforming single-view approaches [11][12] [13][14].
+Currently, multi-view learning has become a consensus in the field of mammography analysis.
+In addition, mammographic images typically possess extremely high resolution, with lesions occupying only a very small area.
+And handling extremely high resolution images is a challenge for traditional networks.
+Such as, in Convolutional Neural Networks (CNNs), layer-by-layer downsampling (like pooling) compresses feature maps, potentially leading to the loss of microcalcifications or subtle structures.
+Accurate localization of these lesions is crucial for diagnosis.
+In the case of Transformers, lesions usually occupy a very small area, and the global attention mechanism of transformers may excessively focus on irrelevant background, introducing noise.
+To address these limitations, several methods have been proposed to integrate global and local information to mitigate the loss of local information.
+The GMIC (Globally-Aware Multiple Instance Classifier) introduces an innovative framework that integrates global and local information extracted by ResNet.
+This concept is akin to multiple instance learning (MIL), where the selection of local information is determined by CAM to identify and extract several local regions of the image [15].
+This method dynamically identifies more important regions for patch-based information extraction using a Saliency Map, then segments the original image to obtain patches.
+It is a weakly supervised approach that requires only image-level labels for patch-level lesion localization.
+This weakly supervised approach effectively reduces the model’s reliance on high-quality annotated data.
+The relevant approach introduces multi-view learning based on GMIC, achieving performance improvement by integrating multi-view feature information through pooling [16].
+These methods are still based on Transformer and CNN, thus the limitations of the traditional paradigms in handling high-resolution images remain unaddressed.
+Therefore, we opted for the COC network to process mammography images [17].
+When using COC (without prior medical image training) for inference, we observed that the network inherently demonstrated a distinct intrinsic clustering capability for breast tissue and lesions, which is visually evident in Figure 1.
+The clustering results may be more easily correlated with anatomical or pathological features and align with clinical needs.
+Figure 1: Context Clustering Visualization Diagram.
+Figures a to d show that the left half of each image shows the original mammogram with annotated suspicious lesions, while the right half presents Contextual clustering visualization, akin to a CNN heatmap and a VIT attention map, with the suspicious lesion locations also outlined.
+This figure clearly shows that the Context clustering approach effectively identifies and groups suspicious lesion areas in mammography.
+Furthermore, these methods of feature fusion across different scales have not effectively utilized the information at each scale.
+For instance, traditional MV methods only employ global information.
+Conventional MIL methods, typically applied to pathological images, can only extract local information, and in GMIC framework methods, the global information extraction network extracts feature information from the entire image, but the feature information that play a role are actually few (due to the characteristic that mammography lesion areas are sparse), resulting in significant loss in the final pooled feature information, and the feature information from the global network are often wasted.
+And, in the current model, the local information is patch-based, lacking connections with other parts.
+It is merely isolated, patch-based information, which inevitably leads to the loss of global characteristics.
+TABLE I: Comparison of Model Structures. There A is Multi-view structure, B is with global information, C is with patch-based local information, D is with feature-based local information.
+Model A B C D
+DsMIL[18]
+AbMIL[19]
+TransMIL[20]
+SV Res[9]
+SV SwinT[21]
+GMIC[15]
+MV Res[9]
+MaMVT[10]
+MV GMIC[16]
+Mammo-Clustering(ours)
+To address the limitations of such feature loss, we propose a three-level information fusion architecture, dividing the features when processing mammography images into global information, feature-based local information, and patch-based local information.
+We believe these three levels of information can maximize the utilization of mammography data.
+Therefore, we propose a 3-level global-local information fusion mechanism, which can be called the Tri-level Information Fusion Framework (TIFF).
+We will compared to previous methods (MV, MIL, GMIC), the differences are shown in the table I.
+We focus on the limitation that the resolution of mammography is too high and where lesions occupy only a small portion of the image, and propose Weakly Supervised Multi-view Tri-level Information Fusion Context Clustering Network.
+Our primary contributions include:
   • $\bullet$
-    Our work is the first to apply a non-CNN and non-attention mechanism image feature extraction method, namely Context Clustering, to the early screening task of mammography for breast cancer.
+    This work introduces a novel non-CNN, non-attention-based feature extraction method using Context clustering for early breast cancer screening in mammography.
   • $\bullet$
-    Our work achieves state-of-the-art accuracy in the early screening task of breast cancer mammography while maintaining the lowest number of parameters.
+    We propose a fusion mechanism named Tri-level Information Fusion Framework (TIFF) that integrates global, feature-based local, and patch-based local information, with enhanced focus on local details.
   • $\bullet$
-    We propose a novel fusion mechanism that integrates global information, feature-based local information, and patch-based local information, placing greater emphasis on local details compared to previous methods.
+    Our method achieves state-of-the-art accuracy with the lowest parameter count among comparable techniques, ensuring efficiency.
 II Method
-Our research aims to develop a rapid and reliable breast cancer early screening system to alleviate the burden on medical professionals and provide screening opportunities for women’s health in underdeveloped regions.
-In this section, we describe the weakly supervised multi-instance multi-view network architecture we propose.
-In the first subsection, we formulate the steps involved in early breast cancer screening through mammography.
-In the second subsection, we provide a detailed description of the framework we propose.
-In the third subsection, we provide a detailed introduction and evaluation of the dataset used.
-In the four subsection, we introduce the evaluation metrics used to assess the effectiveness of early breast cancer screening.
+Our research aims to design a rapid and reliable system for early breast cancer screening, with the dual objectives of reducing the workload on medical professionals and expanding access to screening opportunities for women in underdeveloped regions.
+This section details the architecture of our proposed weakly supervised multi-view network.
+In Subsection 1, we summarize the overall workflow of our method for early breast cancer screening in mammography.
+Subsection 2 provides a comprehensive explanation of the key components of our framework, highlighting their individual contributions to the system’s efficacy.
+Finally, in Subsection 3, we introduce the loss functions utilized during training, elaborating on their roles in optimizing the network’s performance and ensuring robust learning.
 II-A Overall Framework
@@
 The proposed model for mammography classification can be formulated as follows:
-For each image $I$ in the given view, we enhance all points into 5-dimensional information points containing color and position data to obtain the set of points $S\in\mathbb{R}^{5\times w\times h}$, where $w\times h$ is equal with the number of points.
+For each image $I$ in the given view, we enhance all points into 5-dimensional information points containing color and position data to obtain the set of points $S\in\mathbb{R}^{5,w\times h}$, where $w\times h$ is equal with the number of points.
 II-A1 Global Information Extraction
-The set $S$ is input to the first point clustering network to obtain global information $F_{g}$ and Saliency-map $I_{map}$:
-$F_{g},I_{map}=f_{\textbf{global}}(S)$ (1)
-Where $f_{\textbf{global}}$ represents the point clustering network for global information extraction.
-$F_{g}$ is extracted global information from global network.
-The saliency-map extracted by the global network, denoted as $I_{map}$.
-$P=f_{roi}(I_{map})$
-The $P$ is a set of location information, representing the positions of n example patches selected by the ROI selection function $f_{roi}$:
+The set $S$ is input to the first Context-Clustering network to obtain global information $F_{g}$:
+$F_{g}=f_{global}(S)$ (1)
+where $f_{global}$ represents the Context-Clustering network for global information extraction.
+The Saliency Map is obtained by processing the feature information $F_{g}$ extracted by global network through a feedforward network, denoted as $I_{map}$.
+And based on the Saliency map, the ROI selection module outputs a set of location information $P$.
+This selection is independent of dataset annotations, and the positions of the selected n patches can be expressed as:
 $P=\{p_{1},p_{2},\ldots,p_{n}\}$
-where $p_{n}$ is a coordinate representing a position, written as $(x_{n},y_{n})$.
-With $p_{n}$, we can extract n patches $\tilde{I}$ from the original image $I_{i}$ and extract n feature-based local information $F_{fl}$ from the global feature information $F_{g}$.
-$F_{fl}^{n}=F_{g}(x_{n},y_{n})$ (2)
-where $F_{fl}^{n}$ is means n-th feature-based local information.
-II-A2 Local Information Extraction
-Each selected patch $\tilde{I}_{n}$ is treated as a new image, re-enhanced based on each point to obtain its five dimensional point set $\tilde{S}_{n}$.
-And processed through a second point clustering network $f_{\textbf{local}}$ to obtain patch-based local information $F_{pl}^{n}$ :
-$F_{pl}^{n}=f_{\textbf{local}}(\tilde{S}_{n})$ (3)
-where $f_{\textbf{local}}$ represents the point clustering network for local information extraction, similar to global clustering network $f_{\textbf{global}}$.
+where $p_{n}$ is a coordinate representing a position, written as $(x_{n},y_{n})$. The value of n is manually set.
+With $P$, we can extract n patches $\tilde{I}$ from the original image $I$ and extract n feature-based local information $F^{n}_{fl}$ from the global information $F_{g}$.
+For the specified number of patches n, we recommend setting it to 4.
+This is because when the number of suspicious lesion locations exceeds this value, some lesions may be missed, thereby affecting the screening performance of the model.
+However, when the number of patches is greater than the number of suspicious lesion locations, the attention mechanism dynamically adjusts the weights of these patches, allowing the model to focus more on patches that truly contain lesion locations.
+Therefore, in this case, the size does not impact the model’s performance.
+II-A2 Patch-based Local Information Extraction
+Each selected patch $\tilde{I}_{i}$ is treated as a new image, re-enhanced based on each point to obtain its five dimensional point set $\tilde{S}_{i}$.
+And processed through a second Context-Clustering network to obtain patch-based local information $F_{pl}^{i}$, where $i\in\{1,2,...,n\}$.
 II-A3 Information Fusion and Attention Mechanism
-The local information $F_{l}$ from all patches is fused with all feature-based local information and all patch-based local information :
-$F_{l}=F_{fl}\oplus F_{pl}$ (4)
-where Operation $\oplus$ overlays two features.
-after that processed through an attention mechanism to enhance relevant features:
-$F_{a}=f_{attention}(F_{l})$ (5)
-Then, the attention-enhanced information is fused with the original global information, resulting in multi-instance fusion information from single-view:
-$F_{f}=f_{\textbf{fuse}}(F_{g},F_{a})$ (6)
+The local information $F_{l}$ from all patches is generated by fusing all feature-based local information and all patch-based local information:
+$F_{l}=f_{atten}(F_{fl}\oplus F_{pl})=f_{atten}(\sum_{i=0}^{n}F_{g}(x_{i},y_{i})\oplus\sum_{i=0}^{n}f_{local}(\tilde{S}_{i}))$ (2)
+The operation $\oplus$ represents the fusion operate of the two types of information, after that processed through an attention mechanism to enhance relevant features:
+Then, the local information $F_{l}$ is fused with the original global information $F_{g}$, resulting in multi-instance fusion information $F_{f}$ from single-view.
 II-A4 View Fusion and Classification
@@
 This information is then integrated for multi-view fusion, which is used for the final binary classification, resulting in the early screening model’s output:
-$F_{fusion}={f}^{\prime}_{fuse}(F_{f}^{\text{lcc}},F_{f}^{\text{lmlo}},F_{f}^{\text{rcc}},F_{f}^{\text{rmlo}})$ (7)
+$F_{fusion}={f}_{fuse}(F_{f}^{\text{lcc}},F_{f}^{\text{lmlo}},F_{f}^{\text{rcc}},F_{f}^{\text{rmlo}})$ (3)
 where $F_{f}^{\text{lcc}}$ represents the fusion feature of LCC images, other similar situations.
-${f}^{\prime}_{fuse}$ is another fusion structure different from $f_{\textbf{fuse}}$.
-In addition to integrating features $F_{fusion}$, ${f}^{\prime}_{fuse}$ also merges $F_{g}$ and $F_{l}$ from different views.
-$F_{global}={f}^{\prime}_{fuse}(F_{g}^{\text{lcc}},F_{g}^{\text{lmlo}},F_{g}^{\text{rcc}},F_{g}^{\text{rmlo}})$
-$F_{local}={f}^{\prime}_{fuse}(F_{l}^{\text{lcc}},F_{l}^{\text{lmlo}},F_{l}^{\text{rcc}},F_{l}^{\text{rmlo}})$
-Finally, $F_{fusion}$ performs the final classification through $f_{cls}$.
-$\text{res}=f_{cls}(F_{fusion})$ (8)
+${f}_{fuse}$ is a fusion structure.
+In addition to integrating features $F_{fusion}$, ${f}_{fuse}$ also merges $F_{g}$ and $F_{l}$ from different views.
+Finally, $F_{fusion}$ performs the final classification.
 This formulation encapsulates the entire process of the weakly supervised multi-instance multi-view network for mammography classification.
@@
 We enhance each pixel by considering it as a 5-dimensional data point containing color and positional information (r, g, b, x, y).
-After this enhancement, the image can be represented as a set of $h\times w$ 5-dimensional data points, with a scale of $(h\times w,5)$.
+After this enhancement, the image can be represented as a set of $h\times w$ 5-dimensional data points, with a scale of $(h\times w,5)$ [17] .
 We can then perform feature extraction through simple clustering.
@@
 Since each point in the set includes color and positional information, this clustering implicitly incorporates spatial and image information.
-II-B2 Context Cluster Block
-We employ context cluster blocks for hierarchical feature extraction, a paradigm similar to convolutional networks.
-At the beginning of each stage, we utilize a point reducer to decrease the number of points, thereby enhancing computational efficiency.
-Subsequently, a series of context cluster blocks are used to extract deep features and adaptively assign aggregated features to each uniformly selected anchor point within the cluster based on similarity, and connect and fuse the nearly points through linear projection.
-Finally, we perform a point-wise averaging operation on the output of the last layer.
-Figure 2: The figure illustrates the clustering visualization following five steps: selecting central anchor points, identifying neighbors for each anchor, calculating features for each anchor, performing similarity analysis based on these anchors, and representing all clusters on the chart.
+II-B2 Context Clustering
+Context Clustering performs deep feature extraction through a hierarchical structure of Context Clustering Blocks, similar to convolutional networks [17].
+A point reducer is placed before each block to reduce the number of points, thereby enhancing computational efficiency.
+Subsequently, aggregated features are adaptively assigned to each uniformly selected anchor point within clusters based on similarity, and neighboring points are connected and fused through linear projection.
+Figure 3: A visual explanation of Context-Clustering.
+This clustering consists of five components: selecting central anchor points, identifying neighbors for each anchor, calculating features for each anchor, performing similarity analysis based on these anchors, and representing all clusters on the chart.
 As illustrated in Figure 3, the input image undergoes point set transformation, and then, in step a, $n$ central anchor points are uniformly selected in the space.
@@
 $P^{x}=\frac{\left(P^{x}+\sum_{q=1}^{k}p_{q}^{x}\right)}{k+1}$
-where $P^{x}$ represents a $x$ dimension of the central anchor point, $x\in\{r,g,b,h,w\}$. And $p_{q}^{x}$ represents the $n$-th neighbor point in the $x$ dimension, $q\in\{0,1,2,\ldots,k\}$.
+where $P^{x}$ denotes the feature of the central anchor point in the $x$ dimension, $x\in\{r,g,b,h,w\}$.
+Meanwhile, $p_{q}^{x}$ represents the feature of the $q$-th neighboring point in the $x$ dimension, $q\in\{0,1,2,\ldots,k\}$.
 After computing the features for all central anchor points, a similarity analysis is conducted between all points in the point set and each central anchor point’s features in step d.
 Each point is assigned to the cluster of the central anchor point with which it has the highest similarity.
-The steps for the similarity analysis is conducted by computing the pairwise cosine similarity matrix between a point and set of central points:
-$f(P_{i},P^{\prime}_{j})=\frac{P_{i}\cdot P^{\prime}_{j}}{|P_{i}||P^{\prime}_{j}|}$
-where $f(P_{i},P^{\prime}_{j})$ is pairwise cosine similarity compute. $P_{i}$ is $i$-th central anchor point, $i\in\{0,1,2,\ldots,n\}$. $P^{\prime}_{j}$ is $j$-th point in the image point set.
+The steps for the similarity analysis is conducted by computing the pairwise cosine similarity matrix $M$ between a point and set of central points:
+$M(P_{i},P^{\prime}_{j})=\frac{P_{i}\cdot P^{\prime}_{j}}{|P_{i}||P^{\prime}_{j}|}$
+where $M(P_{i},P^{\prime}_{j})$ is pairwise cosine similarity matrix. $P_{i}$ is $i$-th central anchor point, $i\in\{0,1,2,\ldots,n\}$. $P^{\prime}_{j}$ is $j$-th point in the image point set.
 Finally, in step e, all clusters are combined, resulting in the desired clustering outcome for the entire image.
-Since each point contains both feature and positional information, the calculation implicitly emphasizes the point’s distance (locality) and feature similarity.
-Each point is then assigned to the most similar center, resulting in c clusters.
-During this process, clusters may contain varying numbers of points.
-In extreme cases, some clusters may contain zero points, rendering them redundant.
-II-B3 ROI Selection Module
-Firstly, the dimensions of the region of interest mapping are required to be consistent with the dimensions of the saliency-map.
-$h_{crop}=h_{patch}\times h_{map}\div h_{I}$
-$w_{crop}=w_{patch}\times w_{map}\div w_{I}$
-where $h_{patch}$ and $w_{patch}$ are user-defined for the patch image size.
-$h_{map}$ and $w_{map}$ are the size of the saliency-map.
-$h_{I}$ and $w_{I}$ are the size of the original image data.
-$h_{crop}$ and $w_{crop}$ are dynamically adjusted based on the original image and saliency map scales.
-Subsequently, the saliency-map is normalized and divided into regions with a height of $h_{crop}$ and a width of $w_{crop}$ regions for greedy ROI search.
+Since each point in the set includes color and position information, this grouping implicitly incorporates spatial and image information [17].
+II-B3 CAM
+Feature visualization is computed by a feedforward network to generate saliency maps.
+This CAM module consists of convolution layers with a kernel size of 1 and is integrated into the Global Network to enable iterative optimization.
+II-B4 ROI Selection Module
+The saliency map is normalized and divided into regions with a height of $h_{crop}$ and a width of $w_{crop}$ regions for greedy ROI search, $h_{crop}$ and $w_{crop}$ are dynamically adjusted based on the original image and saliency map scales.
 In each iteration, the algorithm greedily identifies each region and selects the one with the largest total weight, as determined by average pooling, among all current regions.
@@
 Figure 4 visualizes the patch-based images selected by the ROISelectModule, along with the patches’ positions on the source image and their comparison with the locations of suspicious lesions.
+TABLE II: Performance of each model on two datasets. SV and MV represent Single-View and Multi-Views, respectively. Backbone of the network after "-".
 Vindr-Mammo CBIS-DDSM
 AUC ACC F1 score AUC ACC F1 score Params
-Single-View Res[37] $0.727\pm 0.02$ 0.783 0.619 $0.719\pm 0.02$ 0.591 0.558 1477025
-Single-View Swin-transformer $0.731\pm 0.02$ 0.651 0.594 $0.724\pm 0.02$ 0.601 0.599 14184625
-GMIC[27] $0.793\pm 0.02$ 0.624 0.518 $0.778\pm 0.02$ 0.712 0.705 22487298
-Multi-View res[37] $0.740\pm 0.02$ 0.753 0.567 $0.731\pm 0.02$ 0.676 0.630 6128546
-MaMVT[38] $0.770\pm 0.02$ 0.918 0.647 $0.749\pm 0.02$ 0.649 0.649 30730082
-Multi-View GMIC[52] $0.797\pm 0.02$ 0.637 0.521 $0.781\pm 0.02$ 0.719 0.699 22686871
-Mammo-Clustering(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.694 $\textbf{0.805}\pm 0.02$ 0.709 0.709 9805459
-TABLE I: Performance of each model on two datasets
-II-B4 Attention Module
-In our model, we use two different attention mechanisms to separately fuse multi-instance and multi-view information.
-Similar Multi-instances information fusion: After the global network, the ROI select module to choose $k$ patch-based images, a number set manually.
+DsMIL[18] $0.605\pm 0.02$ 0.730 0.781 $0.697\pm 0.02$ 0.500 0.583 202116
+AbMIL[18] $0.618\pm 0.02$ 0.776 0.825 $0.726\pm 0.02$ 0.571 0.671 396547
+TransMIL[20] $0.631\pm 0.02$ 0.888 0.890 $0.739\pm 0.02$ 0.635 0.637 2544402
+SV Res18 [9] $0.727\pm 0.02$ 0.783 0.821 $0.719\pm 0.02$ 0.646 0.639 1477025
+SV SwinT[21] $0.731\pm 0.02$ 0.651 0.719 $0.724\pm 0.02$ 0.651 0.599 14184625
+GMIC-Res18 [15] $0.793\pm 0.02$ 0.847 0.878 $0.778\pm 0.02$ 0.682 0.680 22487298
+MV Res18 [9] $0.740\pm 0.02$ 0.753 0.796 $0.731\pm 0.02$ 0.676 0.641 6128546
+MaMVT [10] $0.770\pm 0.02$ 0.882 0.867 $0.749\pm 0.02$ 0.649 0.649 30730082
+MV GMIC-Res18 [16] $0.797\pm 0.02$ 0.887 0.879 $0.781\pm 0.02$ 0.699 0.691 22686871
+MV GMIC-SwinT $0.799\pm 0.02$ 0.874 0.854 $0.785\pm 0.02$ 0.694 0.694 28873234
+Mammo-Clustering(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.906 $\textbf{0.805}\pm 0.02$ 0.709 0.709 9805459
+II-B5 Attention Module
+In our model, we employ attention mechanisms in two locations to integrate multi-instance and multi-view information, respectively.
+Approximate Multi-Instances information fusion: Attention is employed here to mitigate the impact of redundant image patches on training.
+After the global network, the ROI select module to choose $n$ patch-based images, a number set manually.
 This implies not all patch-based images carry beneficial information, and some may be redundant.
@@
 The size of $dim$ varies depending on the model.
-First, we use a neural network layer with simple linear transformations $f_{weights}$ and softmax function to compute attention weights $W_{att}$.
-$W_{att}=\text{softmax}(f_{weights}(F_{l}))$
-$F_{a}=F_{l}\odot W_{att}$
+Firstly, we use a neural network layer with a simple linear transformation $f_{weights}$ and a softmax function to compute attention weights.
+Subsequently, the attention weights are multiplied pointwise with the feature representation of the block-based image $F_{l}$ to obtain the final implicit representation $F_{a}$.
+$F_{a}=F_{l}\odot\text{softmax}(f_{weights}(F_{l}))$
 where $\odot$ represents the stationary point multiplication algorithm.
-Subsequently, the attention weights $W_{att}$ are multiplied pointwise with the feature representation of patch-based images $F_{l}$ is performed to obtain the final implicit representation $F_{a}$.
-Multi-view information fusion: In multi-view learning, not all information from each view is necessarily classified as malignant.
+Multi-view information fusion: Attention is employed here to mitigate the impact of redundant image views on training.
+In multi-view learning, not all information from each view is necessarily classified as malignant.
 However, If a single view exhibits malignant characteristics, the instance should be classified as malignant.
@@
 However, in multi-view attention, this attention module processes not only the $F_{a}$ fused by the multi-instance attention module but also $F_{g}$ and $F_{l}$.
-This is because all three features are considered in the loss function for loss calculation.
-II-B5 Embedding Module
+This is because we will perform targeted optimization on different network structures based on the losses obtained from various features.
+II-B6 Embedding Module
 The embedding module in the model primarily aligns feature-based local information $F_{fl}$ with patch-based local information $F_{pl}$ before their integration.
 Here, we employ a trainable MLP to align the scales.
-We designed relevant ablation experiments to verify its effectiveness.
-II-B6 Maxpooling Module
+We designed relevant ablation experiments to verify its effectiveness, the effectiveness can be observed in Table VI.
+II-B7 Maxpooling Module
 We employ max-pooling to fold and align global information $F_{g}$, facilitating better integration with local information fused through the attention module.
@@
 We chose a composite loss function to achieve targeted optimization of different components.
-$L_{total}=\alpha\cdot L_{global}+\beta\cdot L_{local}+\gamma\cdot L_{fusion}+\delta\cdot L_{map}$
+$L=\alpha\cdot L_{global}+\beta\cdot L_{local}+\gamma\cdot L_{fusion}+\delta\cdot L_{map}$
 After the multi-view fusion module, we retain not only the fused information for regression but also intermediate features such as global information, local information, and saliency maps.
-These features are used to compute a composite loss function for precise optimization of each part of the network.And we determine the sensitivity of the loss function to different types of data through component analysis.
+These features are used to compute a composite loss function for precise optimization of each part of the network.
+And we determine the sensitivity of the loss function to different types of data through component analysis.
 $L_{global}$ is calculated using the global information obtained from multi-view fusion and the ground truth values.
@@
 $L_{fusion}$ represents the model’s final classification error, driving the overall model training.
-The weights $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the proportion of each loss, all manually set.
-Figure 3: Figures a to d provide a visualization of patch-based images extracted by the model.
+The weights $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the proportion of each loss, all manually set, Here, we consider $L_{global}$, $L_{local}$, and $L_{fusion}$ to have equal importance, so we recommend setting $\alpha$, $\beta$, and $\gamma$ uniformly to 1.
+Meanwhile, $L_{map}$ represents the weighted average intensity of the saliency map calculated under the L1 norm, resulting in a relatively large value.
+Therefore, we suggest setting $\delta$ to a value less than 0.001.
+Figure 4: Visualization of patch-based images extracted by the model.
 The green box on the mammography indicates the location of the suspicious lesion, while the blue box represents the patch-based images selected by the model.
@@
 Its comprehensive nature and established use in the research community make it a benchmark for evaluating the performance of mammography-based AI models.
+Both Vindr and CBIS-DDSM provide detailed annotations of lesion locations, but such annotation tasks are generally high-cost.
+By employing weakly supervised learning to enable the network to autonomously localize lesion positions, the cost of dataset creation can be significantly reduced.
+TABLE III: The composition of data for the two datasets.
 Vindr-Mammo CBIS-DDSM
@@
 total 4518 481 4999 814 806 1620
-TABLE II: The composition of data for the two datasets.
 III-B Evaluating Indicator
@@
 Here are the definitions and significance of each metric along with their respective formulas:
-III-B1 AUC (Area Under the Curve)
-AUC means the area under the receiver operating characteristic (ROC) curve.
+III-B1 Classification indicators
+  • AUC (Area Under the Curve): AUC means the area under the receiver operating characteristic (ROC) curve.
 The ROC curve uses the true positive rate for mammography benign-malignant classification as the y-axis and the false positive rate as the x-axis.
@@
 […]
-where $TPR(t)$ is the true positive rate at threshold t, and $FPR(t)$ is the false positive rate at threshold t
-III-B2 ACC (Accuracy)
-Accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined. It gives a straightforward measure of how often the classifier is correct.
+    where $TPR$ is the true positive rate, and $FPR$ is the false positive rate.
+  • ACC (Accuracy): Accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined, the corresponding clinical term is “specificity”.
+It gives a straightforward measure of how often the classifier is correct.
     $\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$
     where $TP$, $TN$, $FP$ and $FN$ represent the numbers of true positives, true negatives, false positives, and false negatives, respectively.
-III-B3 F1 Score
-The F1 Score is the weighted average of Precision and Recall. This score takes both false positives and false negatives into account. It is particularly useful when the class distribution is uneven.
+  • F1 Score: The F1 Score is the weighted average of Precision and Recall.
+This score takes both false positives and false negatives into account.
+Given the long-tail issue in the data, we selected the micro F1 score as the evaluation metric, it is particularly useful when the class distribution is uneven.
 […]
-The F1 score reaches its best value at 1 (perfect precision and recall) and worst at 0.
+III-B2 Localization indicators
+  • MDR(Miss Detection Rate): MDR is defined as the percentage of the number of undetected suspicious lesion areas $N_{miss}$ relative to the total number of suspicious lesion areas $N_{gt}$.
+Because, in clinical practice, we are more concerned about lesions being undetected, i.e., false negatives, rather than false positives.
+    $\textit{MDR}=\frac{N_{miss}}{N_{gt}}$
+  • Recall: Recall is a metric used in object detection to evaluate a model’s ability to identify all relevant objects in an image.
+It measures the proportion of actual positive instances (i.e., objects that should be detected) correctly identified by the model.
+In this context, it reflects the model’s capability to detect all existing lesions.
+    $Recall=\frac{TP}{TP+FN}$
 These metrics collectively provide a comprehensive evaluation of the performance of breast cancer screening models, helping to understand their strengths and weaknesses in various aspects of classification.
 III-C Comparative Experiment
-In this study, we evaluated several models on two datasets: Vindr-Mammo and cbis-ddsm. The performance metrics considered were AUC, ACC, and F1 score.
-III-C1 Vindr-Mammo
-For the Vindr-mammo, ours model achieved the highest performance across all metrics, with an AUC of $0.828\pm 0.02$, ACC of $0.919$, and F1 score of $0.694$.
-For single-view task models, the GMIC with a Global-local structure achieves an AUC of 0.793, significantly outperforming ResNet and Swin-Transformer, which had AUCs of $0.727\pm 0.02$ and $0.731\pm 0.02$, respectively.
-For multi-view task models, our model achieves an AUC of 0.828, significantly surpassing other models.
-Multi-View GMIC also showed competitive performance with an AUC of $0.797\pm 0.02$, validating the effectiveness of the Global-local architecture, but its ACC and F1 score were lower at $0.637$ and $0.512$.
-III-C2 cbis-ddsm
-On the cbis-ddsm dataset, the ours model again demonstrated superior performance with an AUC of $0.805\pm 0.02$, ACC of $0.709$, and F1 score of $0.709$.
-Multi-View GMIC, which had an AUC of $0.781\pm 0.02$ and an F1 score of $0.699$, obtained the highest ACC on this dataset.
-In this dataset, the advantages of the Global-Local architecture are more pronounced, with ResNet and Swin-Transformer showing significant disadvantages in AUC.
-III-C3 Model Complexity
+In this study, we evaluated several models on two datasets: Vindr-Mammo and cbis-ddsm. The performance metrics considered were AUC, ACC, and F1 score and so on.
+III-C1 Assessment of classfication
+For the Vindr-mammo, ours model achieved the highest performance across all metrics, with an AUC of $0.828\pm 0.02$, ACC of $0.919$, and F1 score of $0.906$.
+For Single-View task models, the GMIC-Res18 with a Global-local structure achieves an AUC of 0.793, significantly outperforming other Single-View model, which SV Res18 and SV SwinT had AUCs of $0.727\pm 0.02$ and $0.731\pm 0.02$, respectively.
+For Multi-views task models, our model achieves an AUC of 0.828, significantly surpassing other models.
+Multi-views GMIC-SwinT also showed competitive performance with an AUC of $0.799\pm 0.02$, but its ACC and F1 score were lower at $0.874$ and $0.854$.
+On the CBIS-DDSM dataset, the ours model again demonstrated superior performance with an AUC of $0.805\pm 0.02$, ACC of $0.709$, and F1 score of $0.709$, obtained the highest performance across on this dataset.
+Multi-View GMIC-Res18, which had an AUC of $0.781\pm 0.02$ and an F1 score of $0.699$.
+Noticeably, the Single-View MIL method shows promising potential on this dataset, with TransMIL achieving an AUC of $0.739\pm 0.02$.
+This patch-based information approach significantly outperforms methods that focus solely on global information.
+In this this two datasets, the advantages of Context Clustering and Tri-level Information Fusion architecture are more pronounced than other information fusion method, showing significant advantages in AUC.
+III-C2 Model Complexity
 In terms of model complexity, measured by the number of parameters, the ours model had 98.05 million parameters.
@@
 This is efficient compared to the MaMVT with 30.73 million parameters and the Multi-View GMIC with 22.68 million parameters, considering the performance gains achieved.
-III-C4 ROC curve
-Figure 4: Figures a and b compare the ROC curves of our model with other Single-view and Multi-view architectures on the Vindr-mammo dataset.
+III-C3 ROC curve
+Figure 5: Comparison of ROC curves of different models on two public datasets.
+Figures a and b compare the ROC curves of our model with other Single-view and Multi-view architectures on the Vindr-mammo dataset.
 Figures c and d present the ROC curves comparison on the CBIS-DDSM dataset.
@@
 Additionally, the multi-view learning approach enhances model performance.
+III-C4 Assessment of localization
+We compared the performance of three weakly supervised lesion localization models without comparing them to fully supervised models.
+Our aim was to validate the feasibility of training weakly supervised models on datasets without annotated lesion regions, which would significantly reduce the burden of creating mammography datasets.
+TABLE IV: Comparison of Different Weakly Supervised Models in Lesion Localization Tasks.
+MDR Recall
+GMIC-Res18 0.433 0.476
+MV GMIC-Res18 0.336 0.643
+MV GMIC-Res34 0.412 0.479
+MV GMIC-SwinT 0.322 0.650
+Ours 0.294 0.685
+From Table IV, we observe that our model achieves the lowest missed detection rate (MDR) of 0.294 and the lowest recall rate of 0.476, demonstrating its potential.
+All the comparative models employ the same weakly supervised lesion localization approach as our model, enabling lesion location identification using only classification labels, making them suitable for comparison.
+We selected four different models encompassing both CNN and attention mechanism paradigms and conducted evaluation analysis on two metrics: MDR and recall.
+Compared to the second-best model, our approach shows an improvement of approximately 0.03 in both MDR and recall.
+III-C5 Visual comparison
+Figure 6: Comparison of activation map visualizations from different feature extraction paradigms. The green boxes on the original images indicate the annotated true locations of the lesions.
+We visualized the feature activation maps of different feature paradigms, as shown in Figure 6.
+In this comparison, we selected representative models, ResNet and ViT, as exemplars of CNN and attention mechanisms, respectively, to compare with the Context Clustering paradigm.
+It is evident that the activation maps based on CNN perform the worst.
+although they sometimes identify the locations of regions of interest, they exhibit low contrast.
+Activation maps based on attention mechanisms and clustering both achieve relatively clear and accurate localization of regions of interest.
+However, the activation maps based on Context Clustering are cleaner, indicating they are subject to less interference.
 III-D Ablation Experiment
-III-D1 Different Information Fusion Method
-This ablation study aims to demonstrate the effectiveness of our proposed weakly supervised architecture on mammography.
+III-D1 Different Feature Extraction Paradigm
+This ablation study aims to demonstrate the superiority of Context Clustering in feature extraction performance for Mammography through numerical analysis.
+TABLE V: Performance of different feature extraction paradigm on the Vindr-Mammo
 AUC ACC F1 score
-Single-view Coc $0.762\pm 0.02$ 0.711 0.627
-Multi-View Coc $0.783\pm 0.02$ 0.815 0.658
-Ours(mv+gl) $\textbf{0.828}\pm 0.02$ 0.919 0.694
-TABLE III: Performance of different information fusion methods on the Vindr-Mammo
-The table clearly demonstrates the superiority of our architecture, achieving the highest AUC as well as optimal ACC and F1 scores, indicating its balance in mammography tasks.
-$mv$ represents a multi-view learning approach, and $gl$ refers to our proposed weakly supervised framework.
-III-D2 Different Local Information
+SV Res18(CNN-base) [9] $0.727\pm 0.02$ 0.783 0.821
+SV SwinT(Atten-base) [21] $0.731\pm 0.02$ 0.651 0.719
+SV Coc(Clustering-base) [17] $\textbf{0.762}\pm 0.02$ 0.794 0.833
+The table clearly demonstrates the superiority of the Context Clustering architecture, achieving the highest AUC as well as the best ACC and F1 scores in single-view learning, indicating its balance in mammography tasks.
+$SV$ represents a single-view learning approach.
+III-D2 Different Information Fusion Method
 We identified two distinct sources of local information: patch-based local information and feature-based local information.
 Moreover, this feature-based local information has been overlooked in existing work.
-This ablation study aims to verify the effectiveness of the mechanism integrating feature-based local information with patch-based local information.
+The aim of this ablation study is to validate the effectiveness of our Tri level Information Fusion mechanism that combines these two types of local information with global information.
+TABLE VI: Performance of different information fusion method on the Vindr-Mammo
 AUC ACC F1 score
-patch-based local information $0.810\pm 0.02$ 0.873 0.678
-feature-based local information $0.806\pm 0.02$ 0.895 0.659
-ours $\textbf{0.828}\pm 0.02$ 0.919 0.694
-TABLE IV: Performance of different local information on the Vindr-Mammo
-We found that focusing on only one type of local information doesn’t yield better results.
-The AUC for patch-based local information is 0.810, and for feature-based local information, it’s 0.806.
-However, combining both achieved the best result, with an AUC of 0.828.
+Global $0.783\pm 0.02$ 0.815 0.852
+Global + Patch-based local $0.810\pm 0.02$ 0.890 0.891
+Global + Feature-based local $0.806\pm 0.02$ 0.895 0.868
+TIFF(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.906
+We can clearly found that focusing only on one type of local information does not produce the best results, but it is still significantly better than focusing only on Global Information.
+Adding Patch-based local information to Global Information will increase the AUC to 0.810, and adding Feature-based local information to Global Information will increase the AUC to 0.806.
+However, Tri level Information Fusion mechanism combining the three information achieved the best result with an AUC of 0.828.
 IV Discussion
-Figure 5: From figure a to d, the left half of each image shows the original mammogram with annotated suspicious lesions, while the right half presents our contextual clustering visualization, akin to a CNN heatmap and a VIT attention map, with the suspicious lesion locations also outlined.
-This figure clearly shows that our Context Clustering approach effectively identifies and groups suspicious lesion areas in mammography.
-Figure 5 reveals that lesions in mammography predominantly appear in clustered forms.
-The Coc feature extraction paradigm utilizes clustering of point sets based on color and location information, making it adept at identifying the shape and position of such lesions.
-This paradigm offers simplicity, resulting in excellent interpretability and generalizability.
-Therefore, we believe that the Coc network has a strong capability to learn the prior structure of lesions in mammography images.
-Single-View GMIC Multi-View GMIC Ours
-MDR 0.291 0.277 0.177
-TABLE V: The performance of different models in terms of missed detection rate.
-We introduce an additional evaluation metric, the missed detection rate (MDR):
-$\textit{MDR}=\frac{N_{miss}}{N_{gt}}$
-MDR is defined as the percentage of the number of unrecognized suspicious lesion areas $N_{miss}$ relative to the total number of suspicious lesion areas $N_{gt}$.
-Because, in clinical practice, we are more concerned about lesions being undetected, i.e., false negatives, rather than false positives.
-From Table V, we can see that our model achieved the lowest missed detection rate of 0.177, surpassing other model with rates close to 0.1, demonstrating its potential.
-The other two models, like ours, are weakly supervised methods capable of identifying lesion locations using only classification labels, making them suitable for comparison.
+Although our model achieves state-of-the-art performance on two public datasets, the varying performance of MIL across different datasets has drawn our attention.
+We believe the poor performance of MIL on the Vindr-Mammo dataset is due to its long-tail distribution.
+Addressing the impact of long-tail distribution and exploring MIL’s broader performance on well-balanced datasets will be an interesting research direction in the future.
+Exploring ways to better assist clinicians in screening tasks to reduce their workload also represents a more promising research avenue.
 V Conclusions
-In this study, we developed a novel weakly supervised multi-view model for early breast cancer screening using mammography images.
-Unlike conventional feature extraction paradigms such as CNNs and Transformers, our approach employs a context clustering-based method.
-This strategy allows for the integration of feature-based local information with patch-based local information, enhancing the model’s ability to capture nuanced image details.
-Furthermore, by incorporating multi-view mammography image features, our model effectively leverages complementary information from different perspectives.
-This comprehensive approach addresses the limitations of single-view analysis and improves diagnostic accuracy.
-The model’s performance was rigorously evaluated on two publicly available datasets, Vindr-Mammo and CBIS-DDSM, where it achieved state-of-the-art accuracy, demonstrating its potential as a robust tool for early breast cancer detection.
+This study presents a novel, weakly supervised multi-view tri-level information fusion framework for early breast cancer screening using mammography images.
+Unlike conventional feature extraction paradigms such as CNN and ViT, our approach leverages a context-clustering-based methodology.
+This paradigm enhances computational efficiency and facilitates easier association of structural or pathological features, making it suitable for clinical tasks in mammography.
+Additionally, through the tri-level information fusion framework, the proposed model effectively integrates complementary information from different levels, significantly enhancing diagnostic accuracy.
+Comprehensive evaluations conducted on two publicly available datasets, Vindr-Mammo and CBIS-DDSM, showcase the model’s exceptional performance, achieving state-of-the-art accuracy.
+These results underscore the potential of the proposed framework as a robust and scalable solution for early breast cancer detection, offering significant promise for deployment in clinical settings.